Circular No. 6715

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dc IN NGC 7678 The Beijing Astronomical Observatory (BAO) supernova survey (cf.

IAUC 6612, with also A. Esamdin, Urmqi Astronomical Station) reports an apparent supernova in NGC 7678, discovered on unfiltered CCD frames taken with the 0.6-m reflector at BAO on Aug. 5, 8 and 9. The supernova brightened from unfiltered mag 18.3±0.3 on Aug. 5 to 17.5±0.3 on Aug. 8. The supernova’s position is α = 23^{h}28^{m}28^{s}.41, δ = +22^{o}25′23′′.0 (equinox 2000.0), which is offset 8′′.7 east and 5′′.0 north of the nucleus of the galaxy. Unfiltered CCD images of the same field was taken on 1996 Nov. 26 (limiting mag ∼ 19.0), when no star showed at the position of the supernova.

COMET C/1996 M2 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), the discovery by B. McCarty, University of Birmingham, of another comet in a visual examination of archival SOHO-LASCO C3 coronagraph data. No tail was evident, and the comet was not detected by C2. Measurements by D. A. Biesecker and St. Cyr have been reduced by G. V. Williams and are given on MPEC 1997-P06 in detail, the approximate discovery position being:

1996 UT  α_{2000}  δ_{2000}
June 25.730  6^{h}10.^{m}8  +19^{o}43′

Comet C/1996 M2 is evidently another Kreutz sungrazer. Orbital computations by the undersigned, also given on MPEC 1997-P06, show a strong resemblance to the corresponding computation for comet C/1996 M1, observed eight days previously (IAUC 6713, MPEC 1997-P05), indicating the close physical relationship of this pair.

COMET C/1997 D1 (MUELLER) Total visual magnitude estimates: Apr. 7.90 UT, 13.3 (R. J. Bouma,

Groningen, The Netherlands, 0.25-m reflector); May 9.92, 13.3 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 25.89, 13.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

COMET C/1997 J1 (MUELLER) Total visual magnitudes: July 11.93 UT, 13.8 (M. Lehky, Hradec

Kralove, Czech Republic, 0.42-m reflector); Aug. 5.98, 13.7 (Bouma); 7.04, 13.6: (G. Comello, Roden, The Netherlands, 0.30-m reflector).

1997 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden